OpenStudy (anonymous):

At 2:00 P.M., a thermometer reading 80 degrees F is taken outside, where the are temperature is 20 degrees F. At 2:03 P.M., the temperature reading yielded by the thermometer is 42 degrees F. Later, the thermometer is brought inside, where the are is at 80 degrees F. At 2:10 P.M., the reading is 71 degrees F. When was the thermometer brought indoors?

OpenStudy (anonymous):

that's degrees Fahrenheit :)

OpenStudy (anonymous):

i guess we need the rate of change of the thermometer

OpenStudy (anonymous):

if this is like a newton's law of cooling problem we work with the differences in the temperature. the inital difference in the thermometer and the ouside temperature is 60 degrees, and 3 minutes later it is 22 degrees so the temperature can be modelled by \[60\times (\frac{22}{60})^{\frac{t}{3}}\]

OpenStudy (anonymous):

but my guess is you are supposed to write an exponential equation that looks like \[A_0e^{rt}\] yes?

OpenStudy (anonymous):

that is perhaps start with \[22=60e^{3r}\] and solve for r

OpenStudy (anonymous):

we get \[\frac{11}{30}=e^{3r}\] \[3r=\ln(\frac{11}{30})\] \[r=\frac{1}{3}\ln(\frac{11}{30})\]

OpenStudy (anonymous):

my book says \[du/dt = -k \Delta u\] where u is the change in temperature.

OpenStudy (anonymous):

i get r = -.3344 to four decimal places. then

OpenStudy (anonymous):

yeah that is because they want you to think this is a differential equation, but it is simply an exponential model integrate both sides and you get \[u=A_0e^{-kt}\] and i think we just found k

OpenStudy (anonymous):

in other words we know \[-k=.3344\]

OpenStudy (anonymous):

what's \[A _{o}e ^{rt}\]?

OpenStudy (anonymous):

i used r (for rate) instead of -k

OpenStudy (anonymous):

do you know what newton's law of cooling says?

OpenStudy (anonymous):

change in the difference of the ambient temperature and the temperature of heated (or cooled) object decays exponentially.

OpenStudy (anonymous):

the time rate of change of temperature du/dt is proportional to the temperature difference delta u.. yes?

OpenStudy (anonymous):

yes that is essentially the same thing i said, but you are working with the DIFFERENCES in the temperatures

OpenStudy (anonymous):

so initial difference is 80 - 20 = 60

OpenStudy (anonymous):

that is \[A_0\] the initial differnce in temperature, that is, the difference at time t = 0

OpenStudy (anonymous):

can you show me how did you do it step by step? I want to know how you came up with the answer. and yes, you answered it right! :)

OpenStudy (anonymous):

ok sure, but that is not your final answer, that is just \[-k=-.3344\] or \[k=.3344\] you still have to find the time the thermometer was brought in. we can do that second

OpenStudy (anonymous):

what's k btw? XD

OpenStudy (anonymous):

ok lets start at the beginning

OpenStudy (anonymous):

the initial temperature difference is 60 right?

OpenStudy (anonymous):

okay. i'm really sorry for asking simple questions. it's just that i'm really clueless. :S

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and after 3 minutes the temperature differnce is 22 yes?

OpenStudy (anonymous):

*difference

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

ok now forget about differntial equations for a second

OpenStudy (anonymous):

okay...^^

OpenStudy (anonymous):

i know that the temperature difference decays exponentially, that is, decays proportional to the current amount

OpenStudy (anonymous):

okay. then?

OpenStudy (anonymous):

so i can write an equation for the temperature difference as \[A(t)=A_0e^{-kt}\] where t is time , \[A_0\] is the initial temperature difference (when t = 0) and k is the rate of change. that is what you are looking for, k

OpenStudy (anonymous):

we are given two pieces of information. when t = 0 the temp difference is 60 when t = 3 the temperature diffence is 22 this is like saying i have an exponential function \[A(t)=A_0e^{-kt}\] and i know two points, (0,60) and (3,22)

OpenStudy (anonymous):

so i know what \[A_0\] is right away, when t = 0 we know \[A(0)=A_0e^{-k\times 0}=A_0=60\] more simply put, \[A_0\] is the initial value, which is 60

OpenStudy (anonymous):

ohh.. okay. i'm slowly getting it. :D

OpenStudy (anonymous):

lets make sure it is clear what we are after. we want k

OpenStudy (anonymous):

like finding the slope of a line given two points, only in this case the function is exponential not linear, and k is the rate of change

OpenStudy (anonymous):

now we also know one other point , namely (3,22) because we know after 3 minutes the temperature differnence is 22 degrees.

OpenStudy (anonymous):

so we take our as yes unfinished function \[A(t)=60e^{-kt}\] replace t by 3 and set the result equal to 22

OpenStudy (anonymous):

we get \[A(3)=22=60e^{-k\times 3}\] \[22=60e^{-3k}\] and solve this equation for k

OpenStudy (anonymous):

so we take our as yes unfinished function <~ what? XD

OpenStudy (anonymous):

we don't yet know k that is what we need

OpenStudy (anonymous):

like if you have a line and you know the y intercept is (0,60) so you know it looks like \[f(x)=mx+60\] but you don't yet know m that is what i mean by "as yet unfinished" we still need k

OpenStudy (anonymous):

so our last job is to find k, which we do by solving \[22=60e^{-3k}\] is it clear where i got that equation from?

OpenStudy (anonymous):

setting up that equation is really the whole problem. solving it is easy enough

OpenStudy (anonymous):

yep! super clear! thank you so much! I'm gonna practice how you got that equation now and try to solve other problems. :)

OpenStudy (anonymous):

so now we solve via \[\frac{22}{60}=e^{-3k}\] \[-3k=\ln(\frac{11}{30})\] \[k=-\frac{1}{3}\ln(\frac{11}{30})=.3344\] and our equation is now \[A(t)=60e^{-.3344t}\]

OpenStudy (anonymous):

hmm. okay.

OpenStudy (anonymous):

sorry forgot to finish we now bring the thermometer back into the room. at 2:10 the temperature differnce is 9 so we now have \[A(t)=9e^{-.3344t}\] where again A(t) is the temperature difference, but now t is time starting at 2:10

OpenStudy (anonymous):

we want to know when we brought it back in to the room. at that time the temperature differnce is 80 - 42 = 38 so we have to solve \[38=9e^{-.3344t}\] for t this time we get \[\frac{38}{9}=e^{-.3344t}\] \[t=\frac{\ln(\frac{38}{9})}{-.3344}\] \[t=-4.3\] so 4.3 minutes before 2:10